# Wearable Accelerometer Foundation Models for Health via Knowledge Distillation

### v1

@@ 3.2 Cross-modal representational knowledge distillation
After pre-training the PPG encoder in Section 3.1, we distill its representational knowledge to an accelerometry encoder in a dataset of paired PPG-accelerometry segments (Section 4.1). This second stage is also fully unsupervised without requiring any explicit labels. We perform representational knowledge distillation via multi-modal contrastive learning similar to a technique used previously to weakly supervise an image encoder with text (CLIP) (Radford et al., 2021), but here we use it to transfer knowledge. To do this, unlike standard approaches (Radford et al., 2021), we first perform augmentations on both modalities, PPG and accelerometry, using our stochastic augmentation (see Section 3.1), where the augmentations are independently drawn for each modality in a PPG-accelerometry pair. We found that augmentations were crucial for the quality of the embeddings (see Ablation 5.5). The augmented PPG and accelerometry signals are then processed by the PPG teacher encoder (frozen) and accelerometry student encoder, respectively, to get 256-D output embeddings. To calculate the objective, we first down-project the embeddings to 128-D with trainable multi-layer perceptron projection heads (one 1024-D hidden layer) for both the student and teacher encoders. We found separate learnable projection heads to a smaller subspace was necessary to avoid representation collapse. The student encoder is trained to generate embeddings similar to the teacher encoder, where the objective is contrastive and maximizes the mutual information of paired PPG and accelerometry embeddings, while minimizing the mutual information of an accelerometry embedding with other PPG embeddings. For each batch of embeddings $h$ from N positive pairs for student and teacher ($h_{t}$, $h_{s}$), we define multi-modal InfoNCE, where the teacher embeddings are selected as anchors:
$L^{(t\rightarrow s)}_{\texttt{contrastive}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(sim(h_{t}^{i},h_{s}^{i})/\tau)}{\sum_{j=1}^{N}\exp(sim(h_{t}^{i},h_{s}^{j})/\tau)}.$ (1)
Here, $sim(\cdot,\cdot)$ is the cosine similarity function and $L^{(t\rightarrow s)}_{\texttt{contrastive}}$ can be viewed as a $N$-way classification problem ($N=$ batch size), such that $h_{s}^{i}$ is the correct pair to $h_{t}^{i}$ compared to all other potential pairs in the batch $\{h_{s}^{j}|1\leq j\leq N,j\neq i\}$. The final objective is computed as the weighted sum of InfoNCE, from teacher to student and from student to teacher:
@@ 4.2 Evaluation metrics
PPG-accelerometry pair embeddings retrieval: To assess the quality of the PPG-distilled accelerometry embeddings (256-D representations after the encoder), we perform a retrieval experiment to see how well the accelerometry embeddings can retrieve their corresponding matched PPG segment on unseen test data. For a batch of paired PPG and accelerometry segments on held-out test participants, we compute the cosine similarity between each distilled accelerometry embedding and PPG embedding, which generates a ranked list that can be used for retrieval. We evaluate retrieval quality using top-K accuracy for $K=1,3,5$ (i.e., for a given query accelerometry embedding, how often is its paired PPG embedding in the top $K$ most similar embeddings). We also report the mean rank, i.e., the mean position of the true paired PPG segment in the rankings; smaller values are better, with 1 indicating perfect retrieval as the correct segment is always ranked first. As an ablation, we repeated the same retrieval analysis from PPG to accelerometry embeddings.
Linear probing for downstream targets: As our main targets, we perform linear probing for predicting heart rate (HR), and two popular measures of heart rate variability: standard deviation of normal-to-normal intervals (SDNN) and root mean square of successive differences (RMSSD) (Natarajan et al., 2020; Shaffer & Ginsberg, 2017). In addition, we perform linear probing for predicting self-reported age, body mass index (BMI), biological sex, and 46 health targets including health conditions, medication use, and lifestyle habits, for participants in AHMS. Detailed explanation of our linear probing evaluation metrics is explained in Appendix A.2.
@@ 5.1 Representational knowledge distillation unlocks strong accelerometry-PPG embedding alignment
To quantify the quality of the cross-modality embedding alignment, we performed retrieval analysis from accelerometry embeddings to PPG embeddings (Section 4.2), where PPG embeddings are from the PPG teacher encoder (“PPG-MAE”) and accelerometry embeddings are from the distilled accelerometry encoder (“Accel-KD via PPG-MAE”). To do this, we sample 1,000 random participants from the $\sim$17.5K participants in the test set (no overlap with training), each with an average of 19 PPG and accelerometry segments, and we repeat this procedure 100 times to report standard deviations of our metrics (Table 1). We observed near-perfect alignment of the PPG and accelerometry embeddings after distillation, indicated by the results in Table 1: we achieve an average $99.17$ top-1 accuracy and 1.02 mean rank in random batches of $\sim 19K$ candidates, orders of magnitude better than the baseline encoders. We obtain similarly strong results when repeating this experiment but switching the retrieval task to be instead from PPG embeddings to accelerometry embeddings (Appendix Table 8). To rule out that this near perfect retrieval performance was simply achievable with uni-modal encoder embeddings, we repeated the retrieval analysis using the uni-modal accelerometry embeddings from “Accel-MAE” by applying optimal translation, rotation and scaling via Procrustes alignment (Krzanowski, 2000) to make them as close as possible to “PPG-MAE” embeddings, and observed marked difference in the retrieval performance (Table 1 and Appendix Table 8). Overall, very high retrieval performance (e.g., $99.17$ top-1 accuracy) demonstrates the effectiveness of our representational knowledge distillation framework and how well the distilled accelerometry encoder embeddings match with PPG teacher encoder embeddings. Importantly, these results indicate that distilled accelerometry embeddings may achieve improved performance for predicting downstream targets due to their high alignment with the high-fidelity PPG embeddings, which we will investigate in the next two sections.
Table 2: Distilled accelerometry encoders are predictive of age, BMI and biological sex with high accuracy, and better than baseline accelerometry encoders. Age and BMI metrics are reported with mean absolute error and biological sex with ROC AUC.
@@ 5.2 Representational knowledge distillation from PPG improves the quality of accelerometry embeddings
To gain additional insight about the quality of accelerometry embeddings after knowledge distillation, we compared several encoders in terms of their downstream performance: 1) “Accel-MAE”, 2) “Accel-KD via PPG-MAE”, and 3) supervised encoder trained directly on each target from scratch (“Accel-supervised”). In the meantime, to study the label efficiency of these models, we sweep the proportion of available training segments/labels for linear probing and supervised training from 0.1% to 100% while keeping the number of test segments/labels the same. Figure 2a represents the performance of these encoders, as well as the performance of the “PPG-MAE” teacher encoder (see Appendix Tables 9, 10 and 11 for extended numbers). We observed that the distilled accelerometry encoder (“Accel-KD via PPG-MAE”) not only outperformed “Accel-MAE”, but also was better than the supervised encoder in all targets and all label availability regimes, being particularly robust to the amount of training labels (Figure 2a across x axis) as it retained strong performance even with 1000×1000$\times$1000 × smaller-sized labeled data. This robustness to the amount of labeled data is a major motivation behind keeping our knowledge distillation framework fully unsupervised.
@@
Additionally, to show that our framework is not unique to the pre-training of the teacher encoder or architecture of the teacher/student encoders, we repeated the pre-training and downstream performance comparison between: 1) accelerometry encoder trained with contrastive learning (“Accel-CL”), 2) distilled accelerometry encoder (“Accel-KD via PPG-CL”) from a PPG teacher encoder trained with CL (“PPG-CL”). To simultaneously show robustness to the architecture, in these experiments, we also changed the teacher and student architectures to EfficientNet in the main results (Section 3.1), but conclusions remain the same for contrastive learning with Transformer models (Appendix Table 13). We made similar observations with EfficientNet encoders trained via CL: the distilled accelerometry encoder not only outperformed “Accel-CL” encoder, but also was better than the supervised encoder in all targets and all label availability regimes (Figure 2a; see Appendix Tables 9, 10 and 11 for extended numbers). Interestingly, we made the observation that in uni-modal pre-training, “PPG-MAE” was better than “PPG-CL”, while “Accel-CL” was better than “Accel-MAE” (Figure 2a; see Appendix Tables 9, 10 and 11 for extended numbers), which we further discuss in Appendix A.4.
@@
5.3 Distilled accelerometry encoders are significantly more predictive of demographic variables and health targets
@@
We next questioned whether the distilled accelerometry encoders are predictive of other health-related targets that require capturing waveform information as opposed to pulse timing information that may be sufficient for heart rate and heart rate variability. To this end, we evaluated their downstream prediction performance of a wide array of targets including age, biological sex, BMI and 46 binary health targets derived from AHMS self-reported questionnaires (see Appendix Section A.3). We observed that distilled accelerometry encoders are better predictive of the demographic variables (Table 2) and health targets (Appendix Table 12), compared to the baseline uni-modal accelerometry encoders. This indicates the generalizability of the distilled accelerometry encoders to a wide range of tasks and being a foundation model. To the best of our knowledge, this is the first work demonstrating that a single accelerometry encoder is predictive of demographic variables with remarkably high accuracy (Table 2), and is readily predictive of variety of health targets (Appendix Table 12).
@@ 5.4 Cross-modal representational knowledge distillation can also enable model compression
One important aspect of models for wearable devices is that they should be compact in size for on-device inference with a minimal power footprint. This allows more frequent estimation of digital biomarkers throughout the day, and makes them available from resource-constrained wearable devices. Therefore, we questioned whether we can perform model compression during the representational knowledge distillation from PPG to accelerometry. To show this, we reduced the size of accelerometry student encoder to 4 new smaller sizes by shrinking the depth and width of the Transformer backbone during the distillation (Figure 3, and see Table 7 for model sizes). We observed that our distillation framework robustly maintains the information quantified by downstream performance, even in encoders with significantly smaller sizes. Particularly, our small encoder (“S” in Table 7) is still better than the baseline uni-modal accelerometry encoders (“Accel-MAE” and “Accel-CL”), while being 5$\sim$ 5$\times$ 5 × smaller than the default Transformer model (“XL” in Table 7).
@@ 5.5 Additional ablation studies
Multi-modal pre-training of both PPG and accelerometry encoders simultaneously results in significantly reduced performance: As discussed in Section 2, there are prior works for multi-modal pre-training of biosignals where they bind different modality embeddings via multi-modal contrastive learning (Deldari et al., 2022; Thapa et al., 2024), which does not involve pre-training a uni-modal teacher encoder (stage 1 in Figure 1). While, our motivation here is different for we use multi-modal contrastive learning to enable unsupervised representational knowledge distillation, we questioned whether we get the same improvement for the accelerometry encoders when binding PPG and accelerometry embeddings in a multi-modal pre-training setup. Therefore, we did an ablation of multi-modal contrastive learning where both PPG/accelerometry encoders are trainable during training, and to do a fair comparison, we experimented with different $\lambda$ values. For all different $\lambda$ values, we observed that the learned accelerometry encoder had significantly lower quality embeddings (see Table 3) as quantified by a significant drop of performance in all downstream targets ($95\%$, $47\%$ and $35\%$ higher mean absolute error for heart rate, SDNN, and RMSSD, respectively). In addition, to emphasize the importance of freezing the teacher encoder in knowledge distillation, we did the multi-modal training where we initialized the PPG encoder using the pre-trained weights of “PPG-MAE” encoder, and still observed significant degrade in performance (Appendix Table 19). We believe that this is due to the asymmetric amount of information present in the PPG and accelerometry, such that allowing the PPG encoder to update results in more trivial embeddings and degraded performance for the accelerometry encoder. Overall, this observation further demonstrates the importance of two-stage representational knowledge distillation and freezing the teacher encoder to allow maximal knowledge transfer.
Table 3: Ablation on simultaneous PPG-accelerometry multi-modal pre-training as opposed to two-stage representational knowledge distillation.

### v2

@@ 3.2 Cross-modal representational knowledge distillation
After pre-training the PPG encoder in Section 3.1, we distill its embeddings to an accelerometry encoder in a dataset of paired PPG-accelerometry segments (Section 4.1). This second stage is also fully unsupervised without requiring any explicit labels. We perform this representational knowledge distillation using multi-modal contrastive learning similar to a technique used previously to supervise an image encoder with text (CLIP) (Radford et al., 2021), but here we use it to transfer knowledge. To do this, unlike standard approaches (Radford et al., 2021), we first perform augmentations on both modalities, PPG and accelerometry, using our stochastic augmentation (see Section 3.1), where the augmentations are independently drawn for each modality in a PPG-accelerometry pair. We found that augmentations were crucial for the quality of the embeddings (see Ablation 5.5). The augmented PPG and accelerometry signals are then processed by the PPG teacher encoder (frozen) and accelerometry student encoder, respectively, to get 256-D output embeddings. To calculate the objective, we first down-project the embeddings to 128-D with trainable multi-layer perceptron projection heads (one 1024-D hidden layer) for both the student and teacher encoders. We found separate learnable projection heads to a smaller subspace was necessary to avoid representation collapse. The student encoder is trained to generate embeddings similar to the teacher encoder, where the objective is contrastive and maximizes the mutual information of paired PPG and accelerometry embeddings, while minimizing the mutual information of an accelerometry embedding with other PPG embeddings. For each batch of embeddings $h$ from N positive pairs for student and teacher ($h_{t}$, $h_{s}$), we define multi-modal InfoNCE, where the teacher embeddings are selected as anchors: $L^{(t\rightarrow s)}_{\texttt{contrastive}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(sim(h_{t}^{i},h_{s}^{i})/\tau)}{\sum_{j=1}^{N}\exp(sim(h_{t}^{i},h_{s}^{j})/\tau)}$. Here, $sim(\cdot,\cdot)$ is the cosine similarity function and $L^{(t\rightarrow s)}_{\texttt{contrastive}}$ can be viewed as a $N$-way classification problem ($N=$ batch size), such that $h_{s}^{i}$ is the correct pair to $h_{t}^{i}$ compared to all other potential pairs in the batch $\{h_{s}^{j}|1\leq j\leq N,j\neq i\}$. The final objective is computed as the weighted sum of InfoNCE, from teacher to student and from student to teacher:
@@
Figure 1: Overview of our dataset and methods. We use the multi-modal PPG-accelerometry data collected under informed consent from Apple Watch in Apple Heart and Movement Study. We first pre-train the PPG teacher encoder with masked autoencoding, and then distill its embeddings to an accelerometry encoder via cross-modal knowledge distillation. See Sections 3.1, 3.2, 4.1 and 4.2 for more details.
@@ 4.2 Evaluation metrics
Embedding Top-1 Acc. $\uparrow$ Mean Rank $\downarrow$
Accel-KD via PPG-MAE 99.17 (0.23) 1.02 (0.01)
Accel-MAE + Procrustes align. 0.18 (0.03) 2808.86 (68.95)
Chance-level performance 0.01 9551.64
PPG-accelerometry pair embeddings retrieval: To assess the quality of the PPG-distilled accelerometry embeddings (256-D representations after the encoder), we perform a retrieval experiment to see how well the accelerometry embeddings can retrieve their corresponding matched PPG segment on unseen test data. For a batch of paired PPG and accelerometry segments on held-out test participants, we compute the cosine similarity between each distilled accelerometry embedding and PPG embedding, which generates a ranked list that can be used for retrieval. We evaluate retrieval quality using top-1 accuracy for (i.e., for a given query accelerometry embedding, how often is its paired PPG embedding in the most similar embedding). We also report the mean rank, i.e., the mean position of the true paired PPG segment in the rankings; smaller values are better, with 1 indicating perfect retrieval as the correct segment is always ranked first. As an ablation, we repeated the same retrieval analysis from PPG to accelerometry embeddings.
Linear probing for downstream targets: As our main targets, we perform linear probing for predicting heart rate (HR), and two popular measures of heart rate variability: standard deviation of normal-to-normal intervals (SDNN) and root mean square of successive differences (RMSSD). These targets are chosen due to the importance of predicting them frequently thourought the day, and that they are widely used in wearable devices (Natarajan et al., 2020) and are indicative of health status (Shaffer & Ginsberg, 2017), training load in athletes (Plews et al., 2013) and stress levels (Kim et al., 2018). We also perform linear probing for predicting self-reported age, body mass index (BMI), biological sex, and 46 health targets including health conditions, medication use, and lifestyle habits, for participants in AHMS. More details of our linear probing evaluation metrics is in Appendix A.2.
@@ 5.1 Representational knowledge distillation unlocks strong accelerometry-PPG embedding alignment
To quantify the quality of the cross-modality embedding alignment, we performed retrieval analysis from accelerometry embeddings to PPG embeddings (Section 4.2), where PPG embeddings are from the PPG teacher encoder (“PPG-MAE”) and accelerometry embeddings are from the distilled accelerometry encoder (“Accel-KD via PPG-MAE”). To do this, we sample 1,000 random participants from the $\sim$17.5K participants in the test set (no overlap with training), each with an average of 19 PPG and accelerometry segments, and we repeat this procedure 100 times to report standard deviations of our metrics (Table 1). We observed near-perfect alignment of the PPG and accelerometry embeddings after distillation, indicated by the results in Table 1: we achieve an average $99.17$ top-1 accuracy and 1.02 mean rank in random batches of $\sim 19K$ candidates. We obtain similar strong results when repeating this experiment but switching the retrieval task to be instead from PPG embeddings to accelerometry embeddings (Appendix Table 8). To rule out that the near perfect retrieval performance was simply achievable with uni-modal encoder embeddings, we repeated the retrieval analysis using the uni-modal accelerometry embeddings from “Accel-MAE” by applying optimal translation, rotation and scaling via Procrustes alignment (Krzanowski, 2000) to make them as close as possible to “PPG-MAE” embeddings, and observed marked difference in the retrieval performance (Table 1 and Appendix Table 8). Overall, very high retrieval performance (e.g., $99.17$ top-1 accuracy) demonstrates the effectiveness of our representational knowledge distillation framework and how well the distilled accelerometry encoder embeddings match with PPG teacher encoder embeddings. Importantly, these results indicate that distilled accelerometry embeddings may achieve improved performance for predicting downstream targets due to their high alignment with the high-fidelity PPG embeddings, which we will investigate in the next two sections.
@@ 5.2 Representational knowledge distillation from PPG improves the quality of accelerometry embeddings
To gain additional insight about the quality of accelerometry embeddings after knowledge distillation, we compared several encoders in terms of their downstream performance: 1) “Accel-MAE”, 2) “Accel-KD via PPG-MAE”, and 3) supervised encoder trained directly on each target from scratch (“Accel-supervised”). In the meantime, to study the label efficiency of these models, we sweep the proportion of available training segments/labels for linear probing and supervised training from 0.1% to 100% while keeping the number of test segments/labels the same. Figure 2a represents the performance of these encoders, as well as the performance of the “PPG-MAE” teacher encoder (see Appendix Tables 9, 10 and 11 for extended numbers with other metrics). We observed that the distilled accelerometry encoder (“Accel-KD via PPG-MAE”) not only outperformed “Accel-MAE”, but also was better than the supervised encoder in all targets and all label availability regimes, being particularly robust to the amount of training labels (Figure 2a across x axis) as it retained strong performance even with 1000×1000$\times$1000 × smaller-sized labeled data. This robustness to the amount of labeled data is a major motivation behind keeping our knowledge distillation framework fully unsupervised.
@@
Additionally, to show that this observation is not unique to the pre-training of the teacher encoder or architecture of the teacher/student encoders, we repeated the pre-training and downstream performance comparison between: 1) accelerometry encoder trained with contrastive learning (“Accel-CL”), 2) distilled accelerometry encoder (“Accel-KD via PPG-CL”) from a PPG teacher encoder trained with CL (“PPG-CL”). To simultaneously show robustness to the architecture, in these experiments, we also changed the teacher and student architectures to EfficientNet in the main results (Section 3.1), but conclusions remain the same for contrastive learning with Transformer models (Appendix Table 13). We made similar observations with EfficientNet encoders trained via CL: the distilled accelerometry encoder not only outperformed “Accel-CL” encoder, but also was better than the supervised encoder in all targets and all label availability regimes (Figure 2a; see Appendix Tables 9, 10 and 11 for extended numbers with other metrics). Interestingly, we made the observation that in uni-modal pre-training, “PPG-MAE” was better than “PPG-CL”, while “Accel-CL” was better than “Accel-MAE” (Figure 2a, Appendix Tables 9, 10 and 11), which we further discuss in Appendix A.4.
@@
5.3 Distilled accelerometry encoders are predictive of demographic variables and health targets with high accuracy, i.e., they are generalist foundation models
Table 2: Distilled accelerometry encoders are predictive of age, BMI and biological sex with high accuracy, and better than baseline accelerometry encoders. Age and BMI metrics are reported with mean absolute error in years and $\text{kg}/\text{m}^{2}$ respectively, and biological sex is reported with ROC AUC. See Appendix 12 for other health targets.
Encoder Age $\downarrow$ BMI $\downarrow$ Biological Sex $\uparrow$
Accel-KD via PPG-MAE 4.04 2.48 0.99
Accel-KD via PPG-CL 4.74 2.82 0.97
Accel-MAE 7.73 3.84 0.87
Accel-CL 4.96 2.62 0.98
@@
We next questioned whether the distilled accelerometry encoders are predictive of other health-related targets that require capturing waveform information as opposed to pulse timing information that may be sufficient for heart rate and heart rate variability. To this end, we evaluated their downstream prediction performance of a wide array of targets including age, biological sex, BMI and 46 binary health targets derived from AHMS self-reported questionnaires (see Appendix Section A.3). We observed that distilled accelerometry encoders are predictive of the demographic variables (Table 2) and health targets (Appendix Table 12), and their prediction performance is better compared to the baseline uni-modal accelerometry encoders (Table 2 and Appendix Table 12). This indicates the generalizability of the distilled accelerometry encoders to a wide range of tasks and their capability as a foundation model. To the best of our knowledge, this is the first work demonstrating that a single accelerometry encoder is predictive of demographic variables with high accuracy (Table 2), and is readily predictive of variety of health targets (Appendix Table 12).
@@ 5.4 Cross-modal representational knowledge distillation can also enable model compression
One important aspect of models for wearable devices is that they should be compact in size for on-device inference with a minimal power footprint. This allows more frequent estimation of digital biomarkers throughout the day, and makes them available from resource-constrained wearable devices. Therefore, we questioned whether we can perform model compression during the representational knowledge distillation from PPG to accelerometry. To show this, we reduced the size of accelerometry student encoder to 4 new smaller sizes by shrinking the depth and width of the Transformer backbone during the distillation (Appendix Figure 4, and see Table 7 for model sizes). We observed that our distillation framework robustly maintains the information quantified by downstream performance, even in encoders with significantly smaller sizes. Particularly, our small encoder (“S” in Table 7) is still better than the baseline uni-modal accelerometry encoders (“Accel-MAE” and “Accel-CL”), while being 5$\sim$ 5$\times$ 5 × smaller than the default Transformer model (“XL” in Table 7).
@@ 5.5 Additional ablation studies
Multi-modal pre-training of both PPG and accelerometry encoders simultaneously results in significantly reduced performance: As discussed in Section 2, there are prior works for multi-modal pre-training of biosignals where they bind different modality embeddings via multi-modal contrastive learning (Deldari et al., 2022; Thapa et al., 2024), which does not involve pre-training a uni-modal teacher encoder (stage 1 in Figure 1). While, our motivation here is different for we use multi-modal contrastive learning to enable unsupervised representational knowledge distillation, we questioned whether we get the same improvement for the accelerometry encoders when binding PPG and accelerometry embeddings in a multi-modal pre-training setup. Therefore, we did an ablation of multi-modal contrastive learning where both PPG/accelerometry encoders are trainable during training, and to do a fair comparison, we experimented with different $\lambda$ values. For all different $\lambda$ values, we observed that the learned accelerometry encoder had significantly lower quality embeddings (see Table 3 for $\lambda=1$ and Appendix Table 19 for other $\lambda$ values) as quantified by a significant drop of performance in all downstream targets; $95\%$, $47\%$ and $35\%$ higher mean absolute error for heart rate, SDNN, and RMSSD, respectively. In addition, we showed that in the multi-modal pre-training, even if we initialize the PPG encoder using the pre-trained weights of “PPG-MAE” encoder, we still observe significant degrade in performance (Appendix Table 19). We believe that this is due to the asymmetric amount of information present in the PPG and accelerometry, such that allowing the PPG encoder to update results in more trivial embeddings and degraded performance for the accelerometry encoder. Overall, this observation further demonstrates the importance of two-stage representational knowledge distillation and freezing the teacher encoder to allow maximal knowledge transfer.
Table 3: Ablation on simultaneous PPG-accelerometer multi-modal pre-training compared to two-stage representational knowledge distillation. Numbers are reported in mean absolute error, see Appendix Table 19 for additional numbers.
Pre-training framework Eval. $\downarrow$
Simultaneous multi-modal contrastive learning Heart rate (BPM) 2.36
SDNN (ms) 9.68
RMSSD (ms) 11.30
Cross-modal knowledge distillation (ours, Accel-KD via PPG-MAE) Heart rate (BPM) 1.21
SDNN (ms) 6.58
RMSSD (ms) 8.40